OpenStudy (anonymous):

Billy has 1 gallon of paint. He is going to pour it into a paint tray that measures 10 inches wide, 14 inches long, and 4 cm deep. (1 gallon = 231 in3, 1 inch = 2.54 cm) Which of the following scenarios will happen? The paint will not fill the tray by 441 cm3. The paint will not fill the tray by 10.53 in3. The paint will fill the tray exactly. The paint will overfill the tray by 10.53 in3.

OpenStudy (anonymous):

Volume = 10 * 14 * (4/2.54) Then subtract 231.

OpenStudy (anonymous):

@shelbylynn On the last one I helped you with, it would be A. Because it was not filled all the way.

OpenStudy (anonymous):

0.954426151

OpenStudy (anonymous):

How'd you get that?

OpenStudy (anonymous):

10*14*4/2.54=220.47-231= oh may bad its -10.527

OpenStudy (anonymous):

Correct. We know the gallon of paint poured is 231 and the container can only hold 220.47. Theres an overflow of 10.53. Answer would be D. I had this question too.

OpenStudy (anonymous):

thanks but i think its b